An angle measures 115.8° more than the measure of its supplementary angle. What is the measure of each angle?

Found 2 solutions by ikleyn, amoresroy:
Answer by ikleyn(53756)

.
An angle measures 115.8° more than the measure of its supplementary angle.
What is the measure of each angle?
~~~~~~~~~~~~~~~

Let x be the measure of the greater / (larger) angle ("an angle").

Then its supplement measure is  (x-115.8) degrees.


Since the angles are supplementary,

    x + (x-115.8) = 180  degrees.


Simplify and find x

    2x = 180 + 115.8 = 295.8

     x               = 295.8/2 = 147.9 degrees.


ANSWER. The angles are  147.9 degrees and  180-147.9 = 32.1 degrees.

Solved.



Answer by amoresroy(361)

Let x = the measure of the smaller angle
x + 115.8 = the measure of the larger angle
Supplementary angles add up to 180.
So we have,
x + x + 115.8 = 180
2x = 180 - 115.8
x = 64.2/2
x = 32.1 (smaller angle)
For the larger angle, we have
x + 115.8 = 32.1 + 115.8 = 147.9
Answer:
Smaller angle measures 32.1 degrees.
Larger angle measures 147.9 degrees.
Solved.